Tables of the Parallel Sessions PS 1, Monday Morning

Time  MS4 Room 1  MS7 Room 2  CT Room 3  CT Room 4
10:45  Harel, Guershon  Gaubert, Stephane  Dodig, Marija  Costa, Liliana
11:10 Dogan-Dunlap, Hamide van den Driessche, Pauline Corral, Cristina Fernandes, Rosário
11:35  Maracci, Mirko  McEneaney, William  Fošner, Ajda  Stuart, Jeffrey
12:00  Parraguez, Marcela  Zimmermann, Karel  Horn, Roger  Weaver, James
12:25  Trigueros, María  Ponce, Daniela  Stosic, Marko  Vallejo, Ernesto

PS 2, Monday Afternoon

Time  MS8 Room 1  CT Room 2  CT Room 3  CT Room 4
16:55 McDonald, Judith  Machado, Silvia  Gouveia, María  Furuichi, Shigeru
17:20 Šmigoc, Helena  Poole, George  Domínguez, María Elena Bourgeois, Gerald
17:45 Laffey, Thomas  Vargas, Xaab Nop  Rump, Siegfried M.  Soares, Graça
18:10 Eubanks, Sherod Kopparty, Bhaskara Rao  Bru, Rafael  Uchiyama, Mitsuru
18:35 Catral, Minerva  Vieira, Luis  Grudsky, Sergey  Lee, Hosoo

PS 3, Tuesday Morning

Time MS6 Room 1  MS8 Room 2  CT Room 3  CT Room 4
10:35  Bini, Dario  Szyld, Daniel  Arav, Marina  Cravo, Glória
11:00 Fassbender, Heike Morris, DeAnne Pruneda, Rosa E.  Peña, Marta
11:25 Guo, Chun-Hua  Szyld, Daniel  Semrl, Peter  Roca, Alicia
11:50 Iannazzo, Bruno Neumann, Michael  Glebsky, Lev  Carriegos, Miguel
12:15  Klein, Andre  Gassó, Maria T.  Ferrer, Josep

PS 4, Tuesday Afternoon

Time  MS4 Room 1  MS6 Room 2  MS7 Room 3  CT Room 4
16:55 Klasa-Bompoint, Jacqueline Lancaster, Peter Boimond, Jean-Louis  Russo, Maria Rosaria
17:20  Leon, Steven  Meini, Beatrice  Sergeev, Sergey  Bueno Cachadina, M. Isabel
17:45  Tanguay, Denis  Mena, Hermann  Merlet, Glenn  Frank, Martin
18:10  Martin, William  Zúñiga, Juan Carlos  Butkovic, Peter  Salam, Ahmed
18:35  Group discussion  Martínez, José-Javier

PS 5, Wednesday Morning

Time  MS1 Room 1  MS4 Room 2  MS7 Room 3  CT Room 4
10:35 Brualdi, Richard A.  Moura, Ana  Singer, Ivan  Uhlig, Frank
11:00 Hershkowitz, Danny Zandieh, Michelle Loiseau, Jean Jacques Schaffrin, Burkhard
11:25  Kirkland, Steve  Possani, Edgar  Gavalec, Martin  Wu, Pei Yuan
11:50 Pryporova, Olga  Strong, David  Plavka, Jan  Protasov, Vladimir
12:15  Berman, Avi  Group discussion  Schneider, Hans  Seddighin, Morteza

PS 6, Thursday Morning

Time  MS1 Room 1  MS5 Room 2  CT Room 3  CT Room 4  CT Room 5
10:35 Barrett, Wayne  Gugercin, Serkan  Wrobel, Iwona  Sivic, Klemen  Körtesi, Peter
11:00 van der Holst, Hein Van Dooren, Paul Sánchez, Salvador Marovt, Janko  Im, Bokhee
11:25  Grout, Jason  Beattie, Christopher Furtado, Susana Olshevsky, Vadim  Al Zhour, Zeyad
11:50  Shader, Bryan  Feng, Lihong  Sendov, Hristo Palma, Alejandro  Sebeldin, Anatoly
12:15  Deaett, Louis  Freund, Roland  Fonseca, Carlos  Satô, Kenzi  Shahryari, Mohammad

PS 7, Thursday Afternoon

Time MS2 Room 1  MS3 Room 2  CT Room 3  CT Room 4  CT Room 5
16:55 Dhillon, Inderjit  Rust, Bert W.  Day, Jane  Marques, Maria-da-Graça Prokip, Volodymyr
17:20  Cox, Steve  Bardsley, John  Mikkelson, Rana  Dolinar, Gregor  Rakotondrajao, Fanja
17:45 Meerbergen, Karl  Nagy, James  Perdigão, Cecília  Ahn, Eunkyung  Hanaish, Ibrahim
18:10 Plestenjak, Bor  Hnetynková, Iveta  Mitchell, Lon  da Cruz, Henrique F  Kaibah, Hussein
18:35  Renaut, Rosemary A Sinkovic, John  Lee, Gue Myung

PS 8, Friday Morning

Time MS2 Room 1  MS3 Room 2  CT Room 3  CT Room 4
10:35 Olshevsky, Vadim Estatico, Claudio  Andjelic, Milica  Dopazo, Esther
11:00  Bella, Tom  Aricó, Antonio  Martins, Enide  Bengochea, Gabriel
11:25 Li, Chi-Kwong Stefan, Wolfgang  Schaeffer, Elisa  Castro-González, Nieves
11:50 Karow, Michael  Mead, Jodi  Rodríguez, Juan Alberto  Patricio, Pedro
12:15  Moro, Julio  Narayan, Sivaram  Verde-Star, Luis

PS 9, Friday Afternoon

Time  MS1 Room 1  MS5 Room 2  CT Room 3  CT Room 4
15:30  Olesky, Dale  Baur, Ulrike  Cortés, Vanesa  Guzmán, José Ramón
15:55 Vander Meulen, Kevin Damm, Tobias Teixeira Matos, Isabel  Milligan, Thomas
16:20  DeAlba, Luz  Benner, Peter  Goldberger, Assaf  Barría, José
16:45  Tam, Bit-Shun  Wojciechowski, Piotr  Cheng, Wei
17:10  Esen, Özlem